OpenStudy (anonymous):

How do you find dy/dx when ((lnx)+(lny^2)=3)? Thank you!

OpenStudy (anonymous):

i mainly don't understand how to find dy/dx of lny^2, the rest I believe I can do.

OpenStudy (anonymous):

\[\ln x+\ln y^2=3,\ln x+2\ln y=3,differentiate~ with~ respect~ \to~ x\] \[\frac{ 1 }{x }+\frac{ 2 }{ y }\frac{ dy }{dx }=0\] now you can poceed.

OpenStudy (anonymous):

THANK YOU! that was simpler than I thought.